### v1

@@ Abstract
We study the $S$-wave $D\bar{D}^{*}(I=0)$ scattering at four different pion masses $m_{\pi}$ ranging from 250 MeV to 417 MeV from $N_{f}=2$ lattice QCD. A bound state near the $D\bar{D}^{*}$ threshold likely exists even after considering the possible left-hand cut effect due to the one pion exchange interaction. At $m_{\pi}\approx 417$ MeV where the effective range expansion is valid, the compositeness of the bound state is $X\approx 1$ and indicates a predominant $D\bar{D}^{*}$ component. This state may correspond to $X(3872)$. On the other hand, our results of the finite volume energies also hint at the existence of a $1^{++}$ resonance below 4.0 GeV with a width around 60 MeV.
@@
Introduction.— Ever since its discovery in 2003 Choi et al. (2003), $X(3872)$ (aka $\chi_{c1}(3872)$ Workman et al. (2022)) has still been a hot topic in experimental and theoretical studies. Its quantum number is set to be $I^{G}J^{PC}=0^{+}1^{++}$ Aaij et al. (2013). $X(3872)$ has a very tiny width of $\sim 1~{}\mathrm{MeV}$ and a mass almost at the $D^{0}\bar{D}^{0*}$ threshold. Apart from its large decay rate to $D^{0}\bar{D}^{0*}$, $X(3872)$ also decays into $J/\psi\omega$ (I=0) and $J/\psi\rho^{0}$ (I=1) with comparable branching fractions Workman et al. (2022). The exotic properties of $X(3872)$ have spurred intensive and extensive phenomenological investigations. $X(3872)$ is unlikely a pure conventional $2P$ charmonium whose mass is expected to be above 3.9 GeV Godfrey and Isgur (1985); Barnes et al. (2005), while its mass and decays suggest it to be a loosely bound $D\bar{D}$ molecule Tornqvist (2004); Liu et al. (2008, 2009); Li and Zhu (2012); Qiu et al. (2023); Dong et al. (2021) (and possibly a small fraction of the $c\bar{c}$ component Braaten and Kusunoki (2004); Kalashnikova (2005); Barnes and Swanson (2008); Ortega et al. (2010); Li et al. (2009); Baru et al. (2010); Yamaguchi et al. (2020)). It is also interpreted as a compact tetraquark state Maiani et al. (2005); Ebert et al. (2006). See Refs. Brambilla et al. (2011); Chen et al. (2016); Esposito et al. (2017); Guo et al. (2018); Brambilla et al. (2020); Liu et al. (2019) for reviews.
@@
There are also several lattice QCD studies considering the $c\bar{c}-D\bar{D}^{(*)}$ coupled channel effects Bali et al. (2011); Prelovsek and Leskovec (2013), where a strong coupling between $c\bar{c}$ and $D\bar{D}^{(*)}$ is observed in the axial vector channel ($J^{PC}=1^{++}$). Especially the study in Ref. Prelovsek and Leskovec (2013) finds that this coupling in the axial vector channel results in a bound state below the $D\bar{D}^{*}$ threshold, which can be related to $X(3872)$. However, the study in Ref. Prelovsek and Leskovec (2013) does not pay enough attention to the role played by the possible $\chi_{c1}(2P)$ charmonium in the interpretation of energy levels above the $D\bar{D}^{*}$ threshold. On the hadron level, the interaction between $D$ and $\bar{D}^{*}$ should be mediated by light hadrons and, therefore, sensitive to light quark masses. In this work, we will investigate the quark mass dependence of the $D\bar{D}^{*}$ interaction, which may shed light on the essence of the related binding mechanism. We will extract the scattering parameters, such as the scattering lengths and the effective ranges at different pion masses. The quark mass dependence can provide more information to the existing effective field theories describing the $D\bar{D}^{*}$ interaction AlFiky et al. (2006); Albaladejo et al. (2013); Baru et al. (2013); Wang and Wang (2013); Guo et al. (2018). In lattice QCD, the scattering parameters are usually inferred from the finite volume energy levels of the $D\bar{D}^{*}$ system through Lüscher’s finite volume method Lüscher (1986, 1991); Luscher (1991). According to the quantum numbers of $X(3872)$, we will focus on the $S$-wave $D\bar{D}^{*}$ scattering in the $I=0$ channel. This is numerically challenging because light quark annihilation diagrams contribute to the related correlation functions. For this purpose, we adopt the distillation method Peardon et al. (2009) that provides a sophisticated treatment of both the all-to-all quark propagators and a smearing scheme for quark fields.
@@
Numerical details.— We generate $N_{f}=2$ gauge configurations at four pion masses ranging from 250 MeV to 417 MeV on $16^{3}\times 128$ anisotropic lattices with the anisotropy parameter $\xi=a_{s}/a_{t}\approx 5$, where $a_{s}$ and $a_{t}$ are the spatial and the temporal lattice spacings, respectively. We adopt the tadpole improved gauge action Morningstar and Peardon (1997); Chen et al. (2006) for gluons and the tadpole improved anisotropic clover fermion action Zhang and Liu (2001); Su et al. (2006); Meng et al. (2009) for light sea quarks and the valence charm quark. The spatial lattice spacing $a_{s}$ is set to be $a_{s}=0.136(2)~{}\mathrm{fm}$ through the Wilson flow method Lüscher (2010); Borsányi et al. (2012) at the lowest pion mass (the $w_{0}/a_{s}$ is also measured at other pion masses and a mild quark mass dependence is observed with the largest relative deviation being roughly 1-2%). Then we use the measured $\xi$ from the pion dispersion relation to determine $a_{t}$ at each pion mass. The charm quark mass parameter is tuned to give $(m_{\eta_{c}}+3m_{J/\psi})/4=3069$ MeV. Throughout this study, the correlation functions are calculated using the distillation method Peardon et al. (2009). The perambulators for $u,d$ quarks are calculated using $N_{V}^{(\mathrm{l})}=70$ eigenvectors of the gauge covariant Laplacian operators on the lattice, and the perambulators for charm quark are calculated using $N_{V}^{(\mathrm{c})}=120$. The parameters for the gauge ensembles are listed in Table 1.
With respect to the formation of $X(3872)$, we consider the possible coupled channel effects of $\chi_{c1}$ states and the $S$-wave $D\bar{D}^{*}$. In principle, we should consider also the coupling by the $J/\psi\omega$ system, which, however, is observed to almost decouple and is neglected (see Ref. Prelovsek and Leskovec (2013) and also the Supplemental Materials sup ). The interpolating field operators for these states are built in terms of the smeared $u,d,c$ quark fields. The $c\bar{c}$ operators in this Letter takes the spatially extended $c\bar{c}$ version
@@
For $D\bar{D}^{*}$, we only consider meson-meson operators and ignore the tetraquark-like $qc\bar{c}\bar{q}$ diquark-antidiquark structures. Consider the flavor structure of $D\bar{D}^{*}$ states of the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$, the $D\bar{D}^{*}$ operators take the form
@@
with $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$ are used for operators $(\mathcal{O}_{D},\mathcal{O}_{D^{*}})\equiv(\bar{u}(\bar{d})\Gamma_{D}c,\bar{u}(\bar{d})\Gamma_{D^{*}}c)$.
The $S$-wave operator can be obtained by summing over all the possible directions of the relative momentum, namely, $\mathcal{O}_{AB}^{q}\sim\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q})$, where $R$ runs over all the elements of the octahedral group $O$, and $q=|\mathbf{q}|$ is the magnitude of the relative momentum (in units of $2\pi/L$ with $L$ being the spatial size of the lattice). The operators $\mathcal{O}_{4,6}\equiv\mathcal{O}_{D\bar{D}^{*}}^{q=0,1}$ are involved in the practical calculation. Another $\mathcal{O}_{D\bar{D}^{*}}^{q=0}$ operator (labelled as $\mathcal{O}_{5}$) uses the combination $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$. This build-up is similar to that in Ref. Prelovsek and Leskovec (2013).
Figure 1: Energy levels $E_{n}$ obtained from GEVP analyses at different $m_{\pi}$. In each panel, the dashed lines show the non-interacting $D\bar{D}^{*}$ energies $E_{D\bar{D}^{*}}^{q=0}$ (green) and $E_{D\bar{D}^{*}}^{q=1}$ (blue) in Table 1, while the boxes stand for the finite volume energy levels $E_{n}$ with their heights indicating $1\sigma$ statistical uncertainties. In the left column of each panel are the lowest two $E_{n}$ when only $c\bar{c}$ operators are involved in the GEVP analysis, and the right column includes the lowest four energy levels $E_{n}$ from the three $c\bar{c}$ operators ($\mathcal{O}_{1,2,3}$) and the three $D\bar{D}^{*}$ operators ($\mathcal{O}_{4,5,6}$).
The six operators $\{\mathcal{O}_{\alpha},\alpha=1,2,\ldots,6\}$ are used to calculate the correlation matrix $C_{\alpha\beta}(t)=\langle 0|\mathcal{O}_{\alpha}(t)\mathcal{O}^{\dagger}_{\beta}(0)|0\rangle$. The practical calculation considers all the quark diagrams, except for those involving the charm quark annihilation. By solving the generalized eigenvalue problem (GEVP), $C_{\alpha\beta}(t)v_{\beta}^{(n)}(t,t_{0})=\lambda^{(n)}(t,t_{0})C_{\alpha\beta}(t_{0})v_{\beta}^{(n)}(t,t_{0})$ at given $t$ and $t_{0}$, we obtain the optimized correlation function $C^{(n)}(t)=v^{(n)}_{\alpha}v^{(n)}_{\beta}C_{\alpha\beta}(t)$ which is contributed mainly by the $n$-th state. We then use the ratio method Aoki et al. (2007); Bulava et al. (2024a, b) to extract the finite volume energies $E_{2,3,4}$: we perform one-exponential fit to the ratio function
@@
to extract the energy difference $\Delta_{n}=E_{n}-E_{D\bar{D}^{*}}^{q}$, where $C_{D^{(*)}}(t,q)$ is the correlation function of $D^{(*)}$ with a momentum $q=0,1$. In practice, we use $q=0$ for $E_{2,3}$ and $q=1$ for $E_{4}$. We also perform a two-exponential fit to $C^{(n)}(t)$ in a proper time window for a self-consistent check. Although consistent, the energies $E_{2,3,4}$ from the ratio method have smaller errors than those from the two-exponential fit and are therefore used as our final results (See  sup for more details). All the statistical errors are obtained through the jackknife method.
@@
When only the $c\bar{c}$ operators ($\mathcal{O}_{1,2,3}$) are included in the GEVP analysis, for each $m_{\pi}$ we obtain reliably two energy levels, which are shown as yellow and red boxes on the left hand side of each column in Fig. 1, with their heights indicating the statistical errors. The lowest level $E_{1}^{(0)}$ is roughly 3.5 GeV and corresponds necessarily to the conventional charmonium $\chi_{c1}$, while the second level $E_{2}^{(0)}$ is higher than 3.9 GeV and seemingly coincides with the quark model expectation of the $\chi_{c1}(2P)$ mass Godfrey and Isgur (1985); Barnes et al. (2005); Ebert et al. (2003); Li and Chao (2009); Deng et al. (2017); Wang et al. (2019). It should be noted that this energy level does not necessarily correspond to an eigenstate of the lattice Hamiltonian. When the $\mathcal{O}_{D\bar{D}^{*}}$ operators $\mathcal{O}_{4,5,6}$ are added, more energy levels are obtained. The lowest four energy levels (labeled $E_{1},E_{2},E_{3},E_{4}$ from bottom up) are illustrated with colored boxes on the right of each panel in Fig. 1. No sizable change is observed for $E_{1}$ when compared with the lowest level obtained by $c\bar{c}$ operators. The second lowest level $E_{2}$ (green box) appears close to but right below the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (green dashed line), and $E_{4}$ lies right above the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$ (blue dashed line), see Table 1 for details. It is interesting to see that $E_{3}$ resides between but far from $E_{D\bar{D}^{*}}^{q=0}$ and $E_{D\bar{D}^{*}}^{q=1}$.
@@
Figure 2: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at $m_{\pi}=417~{}\mathrm{MeV}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $\{Z_{\alpha}^{(n)},\alpha=1,2,\ldots,6\}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
@@
The energy levels ($E_{1,2,3,4}$) are taken to be the eigenvalues of the lattice Hamiltonian relevant to $0^{+}1^{++}$ charmonium-like systems. The relative couplings of $\mathcal{O}_{\alpha}$ to different states $Z_{\alpha}^{(n)}=|\langle 0|\mathcal{O}_{\alpha}|n\rangle|$ (normalized by the largest value in each panel) are illustrated in Fig. 2. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. Obviously, the $E_{1}$ state of an energy around $3.5$ GeV is unambiguously assigned to be the $\chi_{c1}(1P)$ state. On the other hand, qualitatively in the language of the state superposition, the $E_{2}$ and $E_{3}$ states are due to the mixing of $\chi_{c1}(2P)$ and non-interacting $D\bar{D}^{*}$ state of $q=0$, which pulls down $E_{2}$ below the $D\bar{D}^{*}$ threshold (apart from the possible attractive interaction of $D\bar{D}^{*}$ due to meson exchanges) and pushes $E_{3}$ upwards, as exactly shown in Fig. 1. It is expected that $E_{4}$ be slightly above the non-interacting energy $E_{D\bar{D}^{*}}^{q=1}$ due to the mixing from $\chi_{c1}(2P)$.
Strictly speaking, $E_{2},E_{3},E_{4}$ are the energy levels of the $D\bar{D}^{*}$ scattering states in a finite box, whose physical implication can be inferred from the corresponding $S$-wave $D\bar{D}^{*}$ scattering amplitude that are conventionally derived through Lüscher’s formula Lüscher (1986, 1991); Luscher (1991). For the single channel $D\bar{D}^{*}$ scattering, one defines for each $E_{n}$ a scattering momentum $p_{n}$ through the relation $E_{n}(p_{n})=\sqrt{m_{D}^{2}+p_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+p_{n}^{2}}$. Then the $S$-wave $D\bar{D}^{*}$ scattering phase shift $\delta_{0}(p_{n})$ reads
@@
Table 2 collects the results of $(p^{2},p\cot\delta_{0}(p))$ of $E_{2,3,4}$ at all the four values of $m_{\pi}$, where the statistical errors are obtained by jackknife analyses.
However, there are subtleties in the derivation of $p\cot\delta_{0}$ for $E_{2}$. The dynamics of the $D\bar{D}^{*}$ scattering may includes the interaction from one pion exchange (OPE), which introduces nonanalyticity to $p\cot\delta_{0}(p)$ when $p^{2}$ is below the left-hand cut (lhc) point $(p_{\mathrm{lhc}}^{1\pi})^{2}=((m_{D^{*}}-m_{D})^{2}-m_{\pi}^{2})/4$ Du et al. (2023); Meng et al. (2023); Raposo and Hansen (2023).
@@
Table 2: The $D\bar{D}^{*}$ scattering parameters. The scattering momentum $p^{2}$, the phase shift $p\cot\delta_{0}(p)$ (and $\delta_{0}$) at $E_{2,3,4}$, as well as the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}\approx\frac{1}{4}[(m_{D}-m_{D^{*}})^{2}-m_{\pi}^{2}]$, are given explicitly for all the four values of $m_{\pi}$. The energy shifts $\Delta_{n}$ (derived from Eq. (3)) of $E_{n}$ from the nearby non-interacting energy of $D\bar{D}^{*}$ are also presented. Also collected are the derived scattering length $a_{0}$, the effective range $r_{0}$, and the estimated binding energy $E_{B}$ for the bound state pole. The asterisks in brackets ‘(*)’ indicate that the corresponding values may be problematic when OPE lhc is considered.

### v2

@@ Abstract
We study the $S$-wave $D\bar{D}^{*}(I=0)$ scattering at four different pion masses $m_{\pi}$ ranging from 250 MeV to 417 MeV from $N_{f}=2$ lattice QCD. Three energy levels $E_{2,3,4}$ are extracted at each $m_{\pi}$. The analysis of $E_{2,3}$ using the effective range expansion (ERE) comes out with a shallow bound state below the $D\bar{D}^{*}$ threshold, and the phase shifts at $E_{3,4}$ indicate the possible existence of a resonance near 4.0 GeV. We also perform a joint analysis to $E_{2,3,4}$ through the $K$-matrix parameterization of the scattering amplitude. In this way, we observe a $D\bar{D}^{*}$ bound state whose properties are almost the same as that from the ERE analysis. At each $m_{\pi}$, this joint analysis also results in a resonance pole with a mass slightly above 4.0 GeV and a width around 40-60 MeV, which are compatible with the properties of the newly observed $\chi_{c1}(4010)$ by LHCb. More scrutinized lattice QCD calculations are desired to check the existence of this resonance.
I Introduction
@@
Ever since its discovery in 2003 Choi et al. (2003), $X(3872)$ (aka $\chi_{c1}(3872)$ Workman et al. (2022)) has been a hot topic in experimental and theoretical studies. It has quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$ Aaij et al. (2013), a mass almost at the $D^{0}\bar{D}^{0*}$ threshold, and a very tiny width of ∼$\sim 1~{}\mathrm{MeV}$. It decays mainly into $D^{0}\bar{D}^{0*}$, and has also decay modes $J/\psi\omega$ (I=0) and $J/\psi\rho^{0}$ (I=1) with comparable branching fractions Workman et al. (2022). The exotic properties of $X(3872)$ imply that it is not a pure $\chi_{c1}(2P)$ charmonium whose mass is expected to be above 3.9 GeV Godfrey and Isgur (1985); Barnes et al. (2005). Phenomenological studies also interpret it to be a loosely bound $D\bar{D}^{*}$ molecule Tornqvist (2004); Liu et al. (2008, 2009); Li and Zhu (2012); Qiu et al. (2024); Dong et al. (2021) (and possibly a small fraction of the $c\bar{c}$ component Braaten and Kusunoki (2004); Kalashnikova (2005); Barnes and Swanson (2008); Ortega et al. (2010); Li et al. (2009); Baru et al. (2010); Yamaguchi et al. (2020)), or a compact tetraquark state Maiani et al. (2005); Ebert et al. (2006). See Refs. Brambilla et al. (2011); Chen et al. (2016); Esposito et al. (2017); Guo et al. (2018); Brambilla et al. (2020); Liu et al. (2019) for reviews.
@@
$X(3872)$ relevant lattice QCD studies observe a strong coupling between $c\bar{c}$ and $D\bar{D}^{(*)}$ Bali et al. (2011); Prelovsek and Leskovec (2013); Padmanath et al. (2015). This coupling can result in a bound state corresponding to $X(3872)$ below the $D\bar{D}^{*}$ threshold Prelovsek and Leskovec (2013); Padmanath et al. (2015). However, the study in Ref. Prelovsek and Leskovec (2013) does not pay enough attention to the role played by the possible $\chi_{c1}(2P)$ in interpreting the energy levels above the $D\bar{D}^{*}$ threshold. On the other hand, the $D\bar{D}^{*}$ interaction can be mediated by light hadrons and therefore can be sensitive to light quark masses. So we will revisit the $X(3872)$ relevant $D\bar{D}^{*}$ scattering at different pion masses and explore the quark mass dependence of the $D\bar{D}^{*}$ interaction, which may provide more information to the existing effective field theories describing the $D\bar{D}^{*}$ interaction AlFiky et al. (2006); Albaladejo et al. (2013); Baru et al. (2013); Wang and Wang (2013); Guo et al. (2018). In lattice QCD, the $D\bar{D}^{*}$ scattering amplitude can be derived from the related finite volume energy levels through Lüscher method Lüscher (1986, 1991a, 1991b). For the $X(3872)$ relevant $D\bar{D}^{*}$ scattering, the major numerical challenge is the calculation of the light quark annihilation diagrams that contribute to the related correlation functions. For this purpose, we adopt the distillation method Peardon et al. (2009) that provides a sophisticated treatment of both the all-to-all quark propagators and a smearing scheme for quark fields.
@@
II Numerical details
We generate $N_{f}=2$ gauge configurations at four pion masses ranging from 250 MeV to 417 MeV on $16^{3}\times 128$ anisotropic lattices. The aspect ratio is tuned to be $\xi=a_{s}/a_{t}\approx 5$, where $a_{s}$ and $a_{t}$ are the spatial and temporal lattice spacings, respectively. We adopt the tadpole improved gauge action Morningstar and Peardon (1997); Chen et al. (2006) for gluons and the tadpole improved anisotropic clover fermion action Zhang and Liu (2001); Su et al. (2006); Meng et al. (2009) for light sea quarks and the valence charm quark. The spatial lattice spacing $a_{s}$ is set to be $a_{s}=0.136(2)~{}\mathrm{fm}$ through the Wilson flow method Lüscher (2010); Borsányi et al. (2012). Then the pion dispersion relation along with the measured $\xi$ are used to determine $a_{t}$ at each pion mass. The charm quark mass parameter is tuned to give $(m_{\eta_{c}}+3m_{J/\psi})/4=3069$ MeV. Throughout this study, the correlation functions are calculated using the distillation method Peardon et al. (2009). We use $N_{V}^{(\mathrm{l})}=70$ in calculating the perambulators for $u,d$ quarks, and $N_{V}^{(\mathrm{c})}=120$ in the calculation of charm quark perambulators, where $N_{V}$ is the number of the eigenvectors of the gauge covariant Laplacian operators on the lattice. The parameters for the gauge ensembles are listed in Table 1.
@@ II Numerical details
The $D\bar{D}^{*}$ operators with the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$ take the meson-meson operator type
@@
with $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$. $S$-wave operators are obtained by summing over all the possible directions of the relative momentum, namely, $\mathcal{O}_{AB}^{q}\sim\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q})$, where $R$ runs over all the elements of the octahedral group $O$, and $q=|\mathbf{q}|$ is the magnitude of the relative momentum (in units of $2\pi/L$ with $L$ being the spatial size of the lattice). The operators $\mathcal{O}_{4,6}\equiv\mathcal{O}_{D\bar{D}^{*}}^{q=0,1}$ are involved in the practical calculation. Another $\mathcal{O}_{D\bar{D}^{*}}^{q=0}$ operator (labeled as $\mathcal{O}_{5}$) uses the combination $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$.
II.1 Finite volume energies
The six operators $\{\mathcal{O}_{\alpha},\alpha=1,2,\ldots,6\}$ are used to calculate the correlation matrix $C_{\alpha\beta}(t)=\langle 0|\mathcal{O}_{\alpha}(t)\mathcal{O}^{\dagger}_{\beta}(0)|0\rangle$. The practical calculation considers all the quark diagrams, except for those involving the charm quark annihilation. By solving the generalized eigenvalue problem (GEVP), $C_{\alpha\beta}(t)v_{\beta}^{(n)}(t,t_{0})=\lambda^{(n)}(t,t_{0})C_{\alpha\beta}(t_{0})v_{\beta}^{(n)}(t,t_{0})$ at given $t$ and $t_{0}$, we obtain the optimized correlation function $C^{(n)}(t)=v^{(n)}_{\alpha}v^{(n)}_{\beta}C_{\alpha\beta}(t)$ which is contributed mainly by the $n$-th state. The finite volume energies $E_{2,3,4}$ are extracted through one-exponential fits to the ratio functions Aoki et al. (2007); Bulava et al. (2024a, b)
@@ II.1 Finite volume energies
where $C_{D^{(*)}}(t,q)$ is the correlation function of $D^{(*)}$ with a momentum $q=0,1$, and $\Delta_{n}=E_{n}-E_{D\bar{D}^{*}}^{q}$ is the difference of $E_{n}$ from the nearest non-interacting $D\bar{D}^{*}$ energy. In practice, we use $q=0$ for $E_{2,3}$ and $q=1$ for $E_{4}$. We also perform two-exponential fits to $C^{(n)}(t)$ in proper time windows for a self-consistent check in Appendix C. All the statistical errors are obtained through the jackknife method.
Figure 1: Energy levels $E_{n}$ obtained at different $m_{\pi}$. In each panel, the dashed lines show the non-interacting $D\bar{D}^{*}$ energies $E_{D\bar{D}^{*}}^{q=0}$ (green) and $E_{D\bar{D}^{*}}^{q=1}$ (blue) in Table 1, while the boxes stand for the finite volume energy levels $E_{n}$ with their heights indicating statistical errors. In the left column of each panel are the lowest two $E_{n}$ when only $c\bar{c}$ operators are involved in the GEVP analysis, and the right column includes the lowest four energy levels $E_{n}$ from all the six operators ($\mathcal{O}_{1-6}$). It should be noted that there may exist an additional energy level corresponding to the $D$-wave ($l=2$) $D\bar{D}^{*}$ state near the blue dashed line in the figure. Similar to the treatment in Ref. Prelovsek and Leskovec (2013), we tentatively neglect the mixing from the $D$-wave wave state to the $S$-wave, which is expected to be suppressed by the phase space factor $p^{2l}$ particularly for the low-energy $D\bar{D}^{*}$ scattering Padmanath and Prelovsek (2022). The effect of this kind of mixing needs to be explored by involving more $D\bar{D}^{*}$ operators.
@@
The obtained finite volume energy levels are illustrated in Fig. 1 by colored boxes, whose heights indicate the statistical errors. In each panel labeled by $m_{\pi}$, the yellow and red boxes in the left column are the two energy levels that are derived from the GEVP analysis involving only the $c\bar{c}$ operators ($\mathcal{O}_{1,2,3}$). The lowest level is roughly 3.5 GeV and corresponds to the conventional charmonium $\chi_{c1}$, while the second level is higher than 3.9 GeV and coincides with the quark model expectation of the $\chi_{c1}(2P)$ mass Godfrey and Isgur (1985); Barnes et al. (2005); Ebert et al. (2003); Li and Chao (2009); Deng et al. (2017); Wang et al. (2019). Note that this energy may not be an eigenvalue of the lattice Hamiltonian. When the $\mathcal{O}_{D\bar{D}^{*}}$ operators $\mathcal{O}_{4,5,6}$ are added, more energy levels are obtained. The colored boxes in the right-hand side of each column in Fig. 1 show the lowest four energy levels (labeled $E_{1},E_{2},E_{3},E_{4}$ from bottom up). $E_{1}$ is almost the same as the lowest level obtained by $c\bar{c}$ operators. The second lowest level $E_{2}$ (green box) appears close to but right below the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (green dashed line), and $E_{4}$ lies right above the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$ (blue dashed line) (see Table 1 for the exact values). It is interesting to see that $E_{3}$ resides in the middle of $E_{D\bar{D}^{*}}^{q=0}$ and $E_{D\bar{D}^{*}}^{q=1}$.
@@
Figure 2: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at $m_{\pi}=417~{}\mathrm{MeV}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $Z_{\alpha}^{(n)}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
@@
$E_{1,2,3,4}$ are taken to be the eigenvalues of the lattice Hamiltonian relevant to $0^{+}1^{++}$ charmonium-like systems. The relative couplings of $\mathcal{O}_{\alpha}$ to different states $Z_{\alpha}^{(n)}=|\langle 0|\mathcal{O}_{\alpha}|n\rangle|$ (normalized by the largest value in each panel) are illustrated in Fig. 2. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. The $E_{1}$ state is naturally the $\chi_{c1}(1P)$ state. The $E_{2}$ and $E_{3}$ states are from the mixing of $\chi_{c1}(2P)$ and non-interacting $D\bar{D}^{*}$ state of $q=0$, which pulls $E_{2}$ downwards below the $D\bar{D}^{*}$ threshold (apart from the possible attractive interaction of $D\bar{D}^{*}$ due to meson exchanges) and pushes $E_{3}$ upwards, as exactly shown in Fig. 1. It is expected that $E_{4}$ is slightly above the non-interacting energy $E_{D\bar{D}^{*}}^{q=1}$ due to the mixing from $\chi_{c1}(2P)$. It should be noted that the $D$-wave $D\bar{D}^{*}$ scattering state can have a $J^{PC}=1^{++}$ quantum number and is degenerate with the $S$-wave state in the non-interacting limit. The $S$-$D$-wave mixing (and also the possible different mixing from $\chi_{c1}(2P)$) may result in two energy levels near $E_{D\bar{D}^{*}}^{q=1}$ (the blue dashed line in Fig. 1). By assuming the contribution from the $D$-wave state is negligible, as indicated in the lattice QCD calculation of the $T_{cc}^{+}(3875)$ relevant $DD^{*}$ scattering Padmanath and Prelovsek (2022); Chen et al. (2022), we use only one operator $\mathcal{O}_{D\bar{D}^{*}}(q=1)$ that resembles the $S$-wave $D\bar{D}^{*}$ scattering state and makes the second energy level indiscernible. The effect of this kind of mixing needs to be explored by involving more $D\bar{D}^{*}$ operators at $q=1$.
Strictly speaking, $E_{2},E_{3},E_{4}$ are the energy levels of the $D\bar{D}^{*}$ scattering states in a finite box and are related to the corresponding scattering amplitude through Lüscher’s formula Lüscher (1986, 1991a, 1991b) for the $S$-wave scattering
@@
where $p$ is the scattering momentum defined through $E_{n}(p_{n})=\sqrt{m_{D}^{2}+p_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+p_{n}^{2}}$ for each $E_{n}$. Table 2 collects the results of $(p^{2},p\cot\delta_{0}(p))$ of $E_{2,3,4}$ at all the four values of $m_{\pi}$, where the statistical errors are obtained by jackknife analyses. However, there are subtleties in the derivation of $p\cot\delta_{0}$ for $E_{2}$. The $D\bar{D}^{*}$ scattering may include the interaction from one pion exchange (OPE), which introduces nonanalyticity to $p\cot\delta_{0}(p)$ when $p^{2}$ is below the left-hand cut (lhc) point $(p_{\mathrm{lhc}}^{1\pi})^{2}=((m_{D^{*}}-m_{D})^{2}-m_{\pi}^{2})/4$ Du et al. (2023); Meng et al. (2024); Raposo and Hansen (2024).
@@
Table 2: $D\bar{D}^{*}$ scattering parameters. The scattering momentum $p^{2}$, the phase shift $p\cot\delta_{0}(p)$ (and $\delta_{0}$) at $E_{2,3,4}$, as well as the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}\approx\frac{1}{4}[(m_{D}-m_{D^{*}})^{2}-m_{\pi}^{2}]$, are given explicitly at four values of $m_{\pi}$. The energy shifts $\Delta_{n}$ (derived from Eq. (2)) of $E_{n}$ from the nearby non-interacting energy of $D\bar{D}^{*}$ are also presented. Also collected are the derived scattering length $a_{0}$, the effective range $r_{0}$ from $E_{2},E_{3}$, and the estimated binding energy $E_{B}$ (and binding momentum $p_{B}=+i|p_{B}|$) for the bound state pole. The from the $K$-matrix parameters and the resulted bound state information are also shown for comparison. The asterisks in brackets ‘(*)’ indicate that the corresponding values may be problematic when OPE lhc is considered.